OpenStudy (anonymous):

Forty-three girls went camping. The girls were able to choose to participate in karate and swimming. Twenty-one girls went swimming, eight participated in both activities, and seven girls did neither. How many girls only did karate?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

add all the girls together then subtract from 43

OpenStudy (ranga):

Draw a Venn Diagram.

OpenStudy (ranga):

 Fill in the numbers.

OpenStudy (ranga):

The answer is not 7.

OpenStudy (anonymous):

whats the answer then?

OpenStudy (ranga):

I am not allowed to give out answers but I can assist you in finding the answer.

OpenStudy (anonymous):

please help meeee

OpenStudy (anonymous):

the answer is 7

OpenStudy (ranga):

In the diagram above first put the number 8 in the area shared by both karate and swimming. Then what number will go in the swimming circle outside the shared area. (Hint: 21 girls went swimming. 8 out of that 21 did BOTH swimming and jkrate).

OpenStudy (anonymous):

omg range it is 7

OpenStudy (ranga):

@peaceout: 8 girls participated in BOTH Karate and Swimming. A TOTAL of 21 girls went swimming. So when you add 21 girls who went swimming and 8 girls who went to swimming and karate, you are DOUBLE COUNTING. Now you are getting 29 girls who went swimming when the total is 21.

OpenStudy (anonymous):

then it would be 6 @ranga

OpenStudy (anonymous):

21+8+8=37 then 43-37=6 @ranga

OpenStudy (ranga):



OpenStudy (ranga):

8 is in BOTH circles. In Swimming Circle, 8 + 13 = 21 (given) Outside both circles (that is neither swimming circle nor karate circle) = 7 Total is 43. So in the Karate circle outside the shared region, the number must be: 43 - 8 - 13 - 7 = 15. So: 15 did Karate ONLY (no swimming) 8 did Karate AND Swimming 13 did Swimming ONLY (no karate) 7 did neither Swimming nor karate. If you add the numbers in the Karate circle it is: 15 + 8 = 23. So 23 girls did Karate.

OpenStudy (anonymous):

ok that makes sense